OpenStudy (anonymous):

A bag contains 3 red marbles, 2 blue marbles and 5 green marbles. What is the probability of selecting a blue marble, replacing it in the bag and then selecting a green marble?

OpenStudy (anonymous):

7/10? @YinYangInsane

OpenStudy (anonymous):

for getting a blue marble there is 2 in 10 chance or 1/5. for selecting a green marble without the blue one gone would be 5 in 10 or 1/2. with the blur marble gone, it would be 5/9

OpenStudy (anonymous):

no direct answers plz

OpenStudy (anonymous):

its not 5/9

OpenStudy (zarkon):

multiply the probabilities (2/10)*(5/10)

OpenStudy (zarkon):

\[=\frac{1}{5}\times\frac{1}{2}=\frac{1}{10}\]